The sum of the digits of a certain two-digit number is 7. Reversing its digits increases the number by 9. what are the two equations?

Found 2 solutions by ewatrrr, lar:
Answer by ewatrrr(10682)


Hi,
Let x and (7-x) represent the original ten and units digits respectively
Question states:
Reversing its digits increases the number by 9
[10x + (7-x)] + 9 = 10(7 -x) + x
Solving for x
9x + 16 = 70 - 9x
18x = 54
x = 3, the original ten's dgit, the original one's digit is 4. (7-3)
CHECKING our Answer
34 + 9 = 43

Answer by lar(3)

Let the digits of the number be x and y.
So the number is xy (note: this does not mean x times y).
x+y=7 and yx=xy+9
In order for yx to equal xy+9, x=y-1.
Now we have the two equations x+y=7 and x=y-1.
x+y=7 -> x=7-y
y-1=7-y
2y=8
y=4
Plug that into one of the original equations: x+4=7
x=3
The number is 34.